OpenStudy (anonymous):

I will give a medal Please check my answer solve the equation. (equation below)

OpenStudy (anonymous):

\[\frac{ w }{ 4 } - 4 = 3\]

OpenStudy (anonymous):

My answer: \[\frac{ w }{ 4 } - 4 + 4 = 3 + 4\] \[\frac{ w }{ 4 } = 7\] \[\frac{ w }{ 4 } \times 4 = 7 \times 4\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

Thanks! @netertareh